OpenStudy (anonymous):

given vectors u = 3i + 9j and v = 3i + 4j, find 5u - 9v in terms of unit vectors. My answer: -12i + 9j?

OpenStudy (whpalmer4):

\[5u -9v = 5(3i+9j) -9(3i+4j) = 15i +45j - 27i -36j = -12i+9j\checkmark\]

OpenStudy (anonymous):

Thank you!